# Model Internal Sleuthing: Finding Lexical Identity and Inflectional Morphology in Modern Language Models

### v1

@@ Abstract
Large transformer-based language models dominate modern NLP, yet our understanding of how they encode linguistic information is rooted in studies of early models like BERT and GPT-2. To better understand today’s language models, we investigate how both classical architectures (BERT, DeBERTa, GPT-2) and contemporary large language models (Pythia, OLMo-2, Gemma-2, Qwen2.5, Llama-3.1) represent lexical identity and inflectional morphology. We train linear and nonlinear classifiers on layer-wise activations to predict word lemmas and inflectional features. We discover that models concentrate lexical information linearly in early layers and increasingly nonlinearly in later layers, while keeping inflectional information uniformly accessible and linearly separable throughout the layers. Further analysis reveals that these models encode inflectional morphology through generalizable abstractions, but rely predominantly on memorization to encode lexical identity. Remarkably, these patterns emerge across all 16 models we test, despite differences in architecture, size, and training regime (including pretrained and instruction-tuned variants). This consistency suggests that, despite substantial advances in LLM technologies, transformer models organize linguistic information in similar ways, indicating that these properties could be fundamental for next token prediction and are learned early during pretraining. Our code is available at https://github.com/ml5885/model_internal_sleuthing
Model Internal Sleuthing: Finding Lexical Identity and Inflectional Morphology in Modern Language Models
[…]
@@ 1 Introduction
Large transformer-based language models (LMs) are widely used for tasks such as text generation, question answering, and code completion (Workshop, 2023; Groeneveld et al., 2024; Grattafiori et al., 2024; Hui et al., 2024). However, how these models internally represent linguistic information remains an active research area. Prior work suggests a hierarchical organization where different layers specialize in capturing distinct levels of linguistic structure, from surface features to syntax and semantics (Jawahar et al., 2019; Tenney et al., 2019; Rogers et al., 2020).
But these studies focus only on first-generation LMs such as BERT and GPT-2 (Devlin et al., 2019; Radford et al., 2019). Since then, language technology has transformed dramatically - today’s models are far larger, trained on much more data, and adapted through extensive post-training procedures (Brown et al., 2020; Groeneveld et al., 2024; Lambert et al., 2025). We ask: how do modern LMs encode linguistic structure, and how do these representations differ from those in earlier models?
Consider the words walk, walked, jump, and jumped. Do language models group words by shared meaning (walk, walked) or by shared grammar (walked, jumped)? More broadly, where and how do LMs encode a word’s lemma (its lexical identity) and its grammatical form (inflectional morphology)?
To answer this, we train classifiers to predict either a word’s lemma or its inflection (e.g., tense, number, degree) from hidden state activations, using annotated sentences from the Universal Dependencies GUM corpus (Zeldes, 2017). We evaluate classifiers on 16 language models spanning diverse architectures and sizes. We find that:
@@
    Lexical information is predominantly encoded in early layers, while inflectional morphology is distributed more uniformly across layers.
@@
    These layer-wise distributions hold across all models, despite large-scale changes in architecture, amount of pretraining, and post-training procedures.
@@
    The effective dimensionality of representations generally declines with layer depth, though in some models the lowest point is in the middle layers - where just one or two dimensions capture nearly all the variance.
2 Model Internal Sleuthing
Figure 1: Overview of our classifier methodology. We extract hidden state activations from each model layer for target words and train two classifiers: one predicting inflection and another predicting lemma. We evaluate three classifier architectures (linear regression, MLP, and random forest) for each task to deduce information accessibility and linear separability across the layers of the language model. Performance differences between linear and non-linear classifiers reveal the encoding structure of linguistic information.
To study how language models encode linguistic information, we train simple probabilistic classifiers on the models’ internal representations or activations. [^1] We use representations, activations, and neurons interchangeably. By observing whether classifiers can recover word-level properties like lemma or inflection from a given representation, we can infer which layers, if any, linguistic information is present, and how directly (e.g., linearly or not). High classifier performance at a given layer suggests that the relevant information is encoded and accessible at that layer. Additionally, comparing performance across layers tells us how lexical and inflectional information is distributed throughout the model.
Figure 1 illustrates our approach. For each layer of each model, we extract hidden state activations for a target word and use them as input features to train specialized classifiers that predict either inflection or lemma. In our experiments, we train three different classifiers: linear regression, two-layer multi-layer perceptron (MLP), and random forest.
2.1 Linear Regression Classifier
Consistent with best practices for probing Hewitt and Liang (2019); Liu et al. (2019), we use simple linear regression classifiers. These models are solved via closed-form ridge regression Hastie et al. (2009), using the equation:
@@ 2.1 Linear Regression Classifier
2.2 Random Forest Classifier
$\hat{y}=\operatorname{argmax}_{j\in\{1,2,...,c\}}s_{j}(x)$ (3)
where $s_{j}(x)$ is the probability that the trained binary classifier assigns to class $j$ for the input $x$ [^2] We use scikit-learn’s implementation (Pedregosa et al., 2011) with the predict_proba() method to obtain probability scores.
2.3 MLP Classifier
To test for more general non-linearity, we train a simple two-layer MLP with ReLU activation, defined as:
@@ 2.3 MLP Classifier
Comparing performance across these three classifier types provides insights into how information is encoded: if linear classifiers perform comparably to more complex models, information is likely linearly encoded; if random forests or MLPs significantly outperform linear models, information is non-linearly encoded (Belinkov and Glass, 2019).
2.4 Training and Evaluation
We stratify all datasets into train, validation and test splits. We train classifiers using the training set, select hyperparameters using the validation set, and evaluate on the held-out test set using accuracy and macro F1. We solve equation 1 in closed-form to identify $W\in\mathbb{R}^{d\times c}$, tune random forests for number of trees & maximum depth, and learn $W_{1}$ and $W_{2}$ for the MLP via stochastic gradient descent using the AdamW optimizer on cross-entropy loss (Kingma and Ba, 2014) [^4] We use weight decay for regularization for MLP models.
Linear Separability Gap
To quantify whether a layer’s information is linearly encoded or not, we define the linear separability gap as the difference in accuracy between non-linear and linear models. A larger gap indicates the signal is buried in non-linear structure, while a small gap suggests it is directly accessible. In practice, this gap is typically between -0.3 and 0.3, though theoretically it can range from -1 to 1.
Model Parameters Pretraining Data
@@ 2.4 Training and Evaluation
BERT Base 110M 12.6B tokens11footnotemark: 1
BERT Large 340M 12.6B tokens11footnotemark: 1
DeBERTa V3 Large 418M 32B tokens11footnotemark: 1
@@
GPT 2 Small 124M 8B tokens11footnotemark: 1
GPT 2 Large 708M 8B tokens11footnotemark: 1
GPT 2 XL 1.5B 8B tokens11footnotemark: 1
Pythia 6.9b 6900M 300B tokens
Pythia 6.9b Tulu 6900M 300B tokens
OLMo 2 7B 7300M 4T tokens
OLMo 2 7B Instruct 7300M 4T tokens
Gemma 2 2B 2610M 2T tokens
Gemma 2 2B Instruct 2610M 2T tokens
Qwen2.5 1.5B 1540M 18T tokens
Qwen2.5 1.5B Instruct 1540M 18T tokens
Llama 3.1 8B 8000M 15T tokens
Llama 3.1 8B Instruct 8000M 15T tokens
Table 1: Overview of models used in this study. 1Converted from GB to tokens using the approximation that 1GB of data is approximately 200M tokens in English (Chang et al., 2024).
Figure 2: Linguistic accuracy of different classifiers across model layers. The top row shows lexeme prediction accuracy, while the bottom row shows inflection prediction accuracy. Columns represent different classifier types: Linear Regression (left), MLP (middle), and Random Forest (right). Each line represents a different model.
@@ 3 Experiments
Using the classifier methodology introduced in § 2, we describe the components of our experimental setup: the dataset, model suite, and procedure for extracting token-level representations.
3.1 Dataset
We use the Universal Dependencies Georgetown University Multilayer (GUM) Corpus in our experiments (Zeldes, 2017). This corpus provides rich annotations for English texts across various genres. We curate a dataset containing 54,816 data points from 8,415 unique sentences, with 8 inflection types, 7,848 unique lemmas, and 11,720 unique word forms.
Each data point corresponds to a target word within its sentence context and includes the full sentence text, the target word’s position in the sentence, its lemma (base form), and its inflection label (grammatical category).
Formally, a data point is a tuple $(S,i,\ell,c)$ where $S=\{w_{0},w_{1},...,w_{n}\}$ is a sentence, $i$ is the index of the target word $w_{i}$ within the sentence, $\ell$ is the lemma of $w_{i}$, and $c$ is the inflection category of $w_{i}$.
The data covers three main syntactic categories: nouns (49.5%), verbs (31.2%), and adjectives (19.4%). The most frequent inflection categories are singular nouns (36.2%), verb base forms (18.4%), and positive adjectives (18.1%), while comparative (0.7%) and superlative (0.5%) forms are less common.
Sentences are relatively concise with an average length of 6.5 words, making them ideal for our analysis as they reduce contextual complexity while providing sufficient context. We split the dataset into training, validation, and test sets with a 70-10-20 ratio using stratified sampling based on inflection labels to ensure balanced representation across all grammatical categories.  [^5] See Appendix B for complete details including tokenization statistics and visualizations
@@ 3.2 Models
We study a diverse set of pretrained transformer language models available via the Hugging Face Transformers library (Wolf et al., 2020). Our selection includes different model architectures (encoder-only, decoder-only), sizes and training regimes. Table 1 lists the models used in this study.
@@ 3.3 Representation Extraction
For each data point $(S,i,\ell,c)$, we tokenize the sentence $S$ using model-specific tokenizers and process it through the model to obtain hidden states from all layers.  [^6] For target words split into multiple subword tokens, we use the representation of the last subword token to represent the entire word (Devlin et al., 2019). These representations form the basis of our layer-wise analysis of how linguistic information is encoded throughout the model.
4 Results

### v4

# Echoes of BERT: Do Modern Language Models Rediscover the Classical NLP Pipeline?

@@ Abstract
Large transformer-based language models dominate modern NLP, yet our understanding of how they encode linguistic information relies primarily on studies of early models like BERT and GPT-2. Building on prior BERTology work, we analyze 25 models spanning classical architectures (BERT, DeBERTa, GPT-2) to modern large language models (Pythia, OLMo-2, Gemma-2, Qwen2.5, Llama-3.1), probing layer-by-layer representations across eight linguistic tasks in English. Consistent with earlier findings, we find that hierarchical organization persists in modern models: early layers capture syntax, middle layers handle semantics and entity-level information, and later layers encode discourse phenomena. We dive deeper, conducting an in-depth multilingual analysis of two specific linguistic properties - lexical identity and inflectional morphology - that help disentangle form from meaning. We find that lexical information concentrates linearly in early layers but becomes increasingly nonlinear deeper in the network, while inflectional information remains linearly accessible throughout all layers. Additional analyses of attention mechanisms, steering vectors, and pretraining checkpoints reveal where this information resides within layers, how it can be functionally manipulated, and how representations evolve during pretraining. Taken together, our findings suggest that, even with substantial advances in LLM technologies, transformer models learn to organize linguistic information in similar ways, regardless of model architecture, size, or training regime, indicating that these properties are important for next token prediction. Our code is available at https://github.com/ml5885/model_internal_sleuthing
@@ 1 Introduction
Large transformer-based language models (LMs) are widely used for tasks such as text generation, question answering, and code completion (Workshop, 2023; Groeneveld et al., 2024; Llama, 2024; Hui et al., 2024) However, how these models internally represent linguistic information remains an active research area. Prior work suggests a hierarchical organization where different layers specialize in capturing distinct levels of linguistic structure, from surface features to syntax and semantics (Jawahar et al., 2019; Tenney et al., 2019a; Rogers et al., 2020).
But these studies focus only on first-generation LMs such as BERT and GPT-2 (Devlin et al., 2019; Radford et al., 2019). Since then, language technology has transformed dramatically - today’s models are far larger, trained on much more data, and adapted through extensive post-training procedures (Brown et al., 2020; Groeneveld et al., 2024; Lambert et al., 2025). We ask: how do modern LMs encode linguistic structure, and how do these representations differ from those in earlier models? In particular, do modern LMs rediscover the classical NLP pipeline suggested by prior probing work (Tenney et al., 2019a)?
In addition to probing for these properties, we perform a focused case study on two linguistic properties: lexical identity and inflectional morphology. Consider the words walk, walked, jump, and jumped. Do language models group words by shared meaning (walk, walked) or by shared grammar (walked, jumped)? More broadly, where and how do LMs encode a word’s lemma (its lexical identity) and its grammatical form (inflectional morphology)?
To answer these questions, we conduct probing experiments for each of these linguistic properties within model representations (hidden state activations). We train classifiers on 25 models spanning different architectures, sizes, and training regimes to determine if they process English linguistic structure in a hierarchical manner. For lexical identity and inflectional morphology, we dig deeper, examining six typologically diverse languages: English, Chinese, German, French, Russian, and Turkish. We also probe attention heads and pretraining checkpoints, find steering vectors, and analyze intrinsic dimensionality, to show where these two linguistic aspects live within layers, how they evolve during pretraining, and how they can be functionally manipulated. We find that:
@@
    Modern LMs consistently discover the classical NLP pipeline: early layers process surface and syntactic information, middle layers emphasize semantic and entity-level information, and later layers capture discourse.
@@
    Lexical information is predominantly encoded in early layers and becomes increasingly non-linear deeper in the network, while inflectional morphology remains linearly accessible across layers and languages.
@@
    Lexical identity and inflectional morphology representations emerge early in pretraining, and are robust across architectures, sizes, and languages. In particular, we find that inflectional features occupy compact, linearly accessible subspaces in the residual stream, enabling effective steering interventions.
2 BERTology: Probe Design and Metrics
Figure 1: Overview of our classifier methodology. We extract hidden state activations from each model layer for target words and train classifiers for token, span and pairwise edge predictions (POS, dependencies, constituents, NER, SRL, SPR, coreference, and relations), as well as word-level lemma and inflection prediction. We compare linear regression, MLP, and random-forest classifiers, compute selectivity using control labels, and summarize where performance emerges with expected layer and center of gravity.
We study how language models encode linguistic information by training simple probabilistic classifiers on layer-wise residual-stream activations. [^1] We use representations, activations, and neurons interchangeably. If a classifier recovers a property (e.g., POS, lemma) from a layer $\ell$, we infer that the information is encoded and accessible at that layer; comparing layers shows how it evolves with depth.
2.1 Probe Design
We follow the edge probing task design of Tenney et al. (2019b). For token tasks (e.g., POS), we use the representation of the last subword token for the target word. For span tasks (e.g., constituents, NER, SRL, SPR), we represent a span by mean-pooling its subword token representations. For pair/edge tasks (e.g., coreference, dependency arcs, relations), we encode a pair $(s_{1},s_{2})$ as $r_{s_{1}};r_{s_{2}};r_{s_{1}}\circ r_{s_{2}};|r_{s_{1}}-r_{s_{2}}|]$, where $r_{s}$ is the token or span representation, $\circ$ is elementwise product, and $|\cdot|$ is absolute value.
We train linear regression and two-layer multi-layer perceptron (MLP) classifiers with cross-entropy loss on top of representations from each individual layer. Unlike Tenney et al. (2019a), who uses scalar mixing weights to fuse information across layers, we do not train probes on combinations of multiple layers. Our goal is analytical rather than predictive: we care about isolating where in the model each property becomes accessible, rather than achieving the highest possible accuracy. Simpler probes trained on individual layers provide a more controlled setting and yield more insightful results about how linguistic information emerges throughout the network. See Appendix˜C for details on classifier hyperparameters and training.
2.1.1 Linear Regression Classifier
Consistent with best practices for probing Hewitt & Liang (2019); Liu et al. (2019), we use simple linear regression classifiers. These models are solved via closed-form ridge regression Hastie et al. (2009), using the equation:
@@ 2.1.1 Linear Regression Classifier
2.1.2 MLP Classifier
To test for more general nonlinearity, we train a simple two-layer MLP with ReLU activation, defined as:
@@ 2.1.2 MLP Classifier
Comparing performance across these classifier types provides insights into how information is encoded: if linear classifiers perform comparably to more complex probes, information is likely linearly encoded; if MLPs significantly outperform linear probes, information is non-linearly encoded (Belinkov & Glass, 2019).
2.2 Metrics
We define several metrics for localizing where information emerges with depth and for quantifying nonlinearity: selectivity, the linear separability gap, and two depth statistics inspired by Tenney et al. (2019a), expected layer and center of gravity.
Control tasks and selectivity.
Probes may simply memorize training data rather than extracting true linguistic information from the representations. To measure this, we follow Hewitt & Liang (2019) and train on randomly permuted labels (control tasks). Selectivity at layer $\ell$ is the difference between real and control accuracy:
$\text{Sel}_{\ell}=\text{Acc}^{\text{real}}_{\ell}-\text{Acc}^{\text{control}}_{\ell}$ (4)
Higher values mean the classifier is extracting true linguistic information rather than memorizing.
Linear separability gap.
We quantify nonlinearity at each layer as the difference in accuracy between a non-linear probe and a linear probe:
$\text{Gap}_{\ell}=\text{Acc}^{\text{nonlin}}_{\ell}-\text{Acc}^{\text{linear}}_{\ell},$ (5)
where positive values indicate useful information is present but not linearly separable; values near zero suggest linear accessibility. The gap lies in [1,1][-1,1] in principle and is typically between 0.3-0.3 and $0.3$ in practice.
Center of gravity.
Let $a_{\ell}$ be the test accuracy using layer $\ell$ for $\ell=0,\dots,L$, and let $b_{\ell}=\max_{j\leq\ell}a_{j}$ be the cumulative (best-so-far) curve. We weight layers by their consolidation relative to the baseline and take an index-weighted average:
$w_{\ell}=\frac{b_{\ell}-b_{0}}{\sum_{k=0}^{L}(b_{k}-b_{0})},\quad\text{CenterOfGravity}=\sum_{\ell=0}^{L}\ell\,w_{\ell}.$ (6)
Higher values indicate that performance is consolidated in higher layers.
Expected layer.
To localize where marginal gains arise, we use the nonnegative increments of the cumulative curve and take their index-weighted average:
$\Delta_{\ell}=\max(b_{\ell}-b_{\ell-1},0)\ ,\quad p_{\ell}=\frac{\Delta_{\ell}}{\sum_{j=1}^{L}\Delta_{j}},\quad\text{ExpectedLayer}=\sum_{\ell=1}^{L}\ell\,p_{\ell}.$ (7)
Unlike center of gravity (which weights consolidated performance), this emphasizes where useful information first becomes available, highlighting the specific layers at which the model begins to encode properties relevant to the task.
Model Parameters Pretraining Data Layers
@@ 2.2 Metrics
BERT Base 110M 12.6B tokens11footnotemark: 1 12
BERT Large 340M 12.6B tokens11footnotemark: 1 24
DeBERTa V3 Large 418M 32B tokens11footnotemark: 1 24
@@
GPT-2-Small 124M 8B tokens11footnotemark: 1 12
GPT-2-Large 708M 8B tokens11footnotemark: 1 36
GPT-2-XL 1.5B 8B tokens11footnotemark: 1 48
Goldfish English 1000mb 124M 200M tokens 12
Goldfish Chinese 1000mb 124M 200M tokens 12
Goldfish German 1000mb 124M 200M tokens 12
Goldfish French 1000mb 124M 200M tokens 12
Goldfish Russian 1000mb 124M 200M tokens 12
Goldfish Turkish 1000mb 124M 200M tokens 12
Pythia-6.9B 6900M 300B tokens 32
Pythia-6.9B Tulu 6900M 300B tokens 32
OLMo-2-7B 7300M 4T tokens 32
OLMo-2-7B-Instruct 7300M 4T tokens 32
Gemma-2-2B 2610M 2T tokens 26
Gemma-2-2B-Instruct 2610M 2T tokens 26
Qwen2.5-1.5B 1540M 18T tokens 28
Qwen2.5-1.5B-Instruct 1540M 18T tokens 28
Qwen2.5-7B 7620M 18T tokens 28
Qwen2.5-7B-Instruct 7620M 18T tokens 28
Llama-3.1-8B 8000M 15T tokens 32
Llama-3.1-8B-Instruct 8000M 15T tokens 32
Encoder-Decoder
mT5-base 580M 1T tokens 12
Table 1: Overview of models used in this study
1Converted from GB to tokens using the approximation that 1GB of data is approximately 200M tokens in English (Chang et al., 2024).
@@ 3 Experiments
Using the methodology introduced in §˜2, we describe the components of our experimental setup: the datasets, model suite, and procedure for extracting token-level representations.
3.1 Datasets
We evaluate the eight tasks used by Tenney et al. (2019a): UD English-GUM (POS, dependencies, named entities, coreference, constituents) (Nivre et al., 2016; Zeldes, 2017), Universal Propositions English-EWT (SRL) (Jindal et al., 2022), SPR1 datasets (PropBank and UD-EWT sources; SPR), and SemEval-2010 Task 8 (relations). We use the same token/span/edge labeling schemes.
For our in-depth analysis of lexical identity and inflectional morphology, we use Universal Dependencies corpora across six languages - English, Chinese, German, French, Russian, Turkish (Nivre et al., 2016). We select GUM for English (Zeldes, 2017), GSD for Chinese/German/French (McDonald et al., 2013; Guillaume et al., 2019), SynTagRus for Russian (Droganova et al., 2018), and IMST for Turkish (Sulubacak et al., 2016) [^3] See Appendix F for complete details including dataset statistics, tokenization information, and visualizations for all languages
@@ 3.2 Models
We study a diverse set of pretrained transformer language models spanning different architectures, sizes, and training regimes. Table˜1 lists all models used in this study (see Table˜7 for the HuggingFace identifiers).
For English, we evaluate all models listed in Table˜1 (excluding the non-English Goldfish models). For the five non-English languages (Chinese, German, French, Russian, Turkish), we focus on models that have explicit multilingual training: the Goldfish monolingual models trained specifically for each target language (Chang et al., 2024), multilingual Qwen2.5 variants that include these languages in their training data, and the multilingual mT5-base model (Xue et al., 2021). This ensures that we evaluate models on languages they were trained on while maintaining sufficient coverage.
@@ 3.3 Representation Extraction
We tokenize inputs with model-specific tokenizers and run a forward pass to collect residual-stream activations from every layer. Token, span, and pair encodings follow §˜2. For words split into multiple subwords, we use the last subword’s representation (Devlin et al., 2019). Models are used as-is (no fine-tuning), and we report results by layer using these frozen activations.
4 The Classical NLP Pipeline
In this section, we present probing results for eight English tasks that span the classical NLP pipeline. We report layer-wise accuracy for linear and MLP probes, summarize where each property emerges with Expected Layer and Center of Gravity, and compare models by correlating their layer-by-task accuracy profiles. All figures we show here are for three representative models; chosen to cover architecture, size and training regime. Full results for all 25 models appear in Appendix˜D.
4.1 Layer-wise accuracy
Figure 2: Left: heatmaps of probe accuracy by layer for three models (BERT-Base, Qwen2.5-1.5B, OLMo-2-1124-7B-Instruct). The top row shows MLP probe accuracy, the bottom row shows linear probe accuracy. Right: Pearson correlations between models are computed by vectorizing each model’s per-layer, per-task accuracy grid. The lower triangular matrix shows MLP probe accuracy correlations, the upper triangular matrix shows linear probe accuracy correlations.
Figure 2 shows accuracy heatmaps for three representative models and correlations between the task accuracy all models. In general, we observe high probe accuracy for tasks such as NER and low accuracy for tasks such as semantic relation extraction. Mid-level semantic tasks such as SRL and SPR peak in middle layers, while later layers capture discourse-level phenomena like Coreference and Relations. To compare results across models, we compute Pearson correlations between the task-by-layer accuracy vectors for each pair of models. We find that encoder-only models cluster together, decoder-only models form a separate cluster, and cross-family correlations are weaker. With linear probes, model size also separates clusters: small/medium decoder-only models (100M-2B) correlate strongly with each other, while 7B+ models form a distinct high-similarity group.
4.2 Summary Statistics: Expected Layer and Center of Gravity
Figure 3: Expected Layer (blue, equation˜7) and Center of Gravity (purple, equation˜6) for the same three models. $\tau_{\text{lin}}$ and $\tau_{\text{MLP}}$ give selectivity (real vs. control accuracy) for linear and MLP probes. Higher $\tau$ indicates probes extract true signal rather than memorizing.
The summary statistics in Figure 3 demonstrate a more nuanced pattern. Consistent with results from Tenney et al. (2019a), we find that for early models like BERT-Base and BERT-Large, the hierarchical structure is pronounced: syntax peaks earliest, semantics and entities in the middle, and discourse phenomena latest. For larger modern models like Qwen2.5-1.5B and OLMo-2-7B-Instruct, information emerges earlier, indicated by Expected Layer values shifting toward shallower layers. These models compress multiple linguistic levels into fewer layers, perhaps reflecting stronger pretraining signals and greater model capacity. However, we still observe a hierarchical progression in center-of-gravity metrics, indicating that later layers still play a role in processing linguistic information. Selectivity scores confirm that the probe performance reflects real structure rather than memorization, especially for mid- and late-stage tasks.
4.3 Do Modern LMs Rediscover the Classical NLP Pipeline?
Overall, our results demonstrate that modern language models consistently rediscover the classical NLP pipeline observed in prior BERTology work. Across architectures, training regimes, and model scales, we find that syntactic information is typically represented in earlier layers, semantic information becomes salient in intermediate layers, and discourse-level features are encoded most strongly in later layers. However, we also find that model capacity affects where in the network this information emerges. Smaller, earlier models such as BERT-Base exhibit the most clearly separated stages, with each linguistic property peaking at a distinct depth. As models become larger and more powerful, the peaks for different linguistic properties shift toward earlier layers, indicating that modern models learn rich representations more quickly and need fewer layers to consolidate linguistic knowledge. These findings suggest that while the hierarchical organization of linguistic information remains a robust property of language model representations, powerful models build useful representations earlier in the network
5 Lexical Identity and Inflectional Morphology
Having established that modern LMs represent the steps of the classical NLP pipeline in a hierarchical manner (§˜4), we now present an in-depth analysis of two token-level properties: lexical identity and inflectional morphology. We employ the same probing setup as in §˜4, but expand our analysis to six typologically diverse languages: English, Chinese, German, French, Russian, and Turkish. Finally, we examine where and how these properties are encoded in model representations, including which layers make them linearly accessible.
5.1 Results